EA011B

PI: Alef
Title: Cyg A

Session: Oct 1996 (session 3)
Frequency: 3.6 cm
Mode: B
Obs time (UT): 305 EU 09:30(31/10)-21:30(31/10)
Stations: EbNtOnShSmUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Huang Xinyong at Tue Nov 5 03:55:30 EST 1996) Minor failures
Clock offset = 0.2 micro sec. Data is invalid from 9:35 to 9:54 and 15:28 to 15:55, because of antenna problems. Weather was overcast.

Onsala (J. conway, A. Polatidis at Tue Nov 5 05:24:17 EST 1996) Success
No known problems. Light rain through most of the experiment.

Westerbork (tony foley at Tue Nov 5 06:28:50 EST 1996) Failure
Total power levels for S and X band very unstable and so impossible to produce LCP from linear polarization. Linear polarization recorded.

OAN\-Yebes (Pablo de Vicente at Tue Nov 5 08:57:52 EST 1996) Minor failures
Tracking problems between 10:04:00 14:09:04. Computer crash between 17:35:00 - 18:02:34

Noto (Corrado Trigilio at Tue Nov 5 11:47:58 EST 1996) Minor failures
Some problem of communication with BBC03 at 17:22 for 7 minuts. Lost 8 minuts for antenna in prelimit from 17:55 to 17:59. Cloudy sky.

Effelsberg (d graham at Wed Nov 6 09:44:07 EST 1996) Minor failures
0935-0948 wrong frequency, 1542-1550 not on srce

Urumqi (Liu Xiang at Tue Nov 19 03:46:55 EST 1996) Success
Clock offset 0.1+/-0.4us.

MPIfR (Walter Alef at Tue Sep 23 17:31:39 METDST 1997) Minor failures
Westerbork was not correlated. Rest O.K. Fringes for: Eb, Sh, Sm, Nt, On, Ur, Wz, Yb.